What is the range of convictions in 1990? In 2009? Which is greater?



Number of Public C 1990 2009 Govt Level No. of Convictions Govt.Level No of Convictions Federal426 Stale Local Federal 83 Stale 79 102 Local 225 257
Answer #1

range = maximum-minimum  
  
for 1990  
range = 583-79=   504
  
for 2009  
range = 426-102=   324
  
so, range of convictions is greater for 1990